# A general approach to modeling environmental mixtures with multivariate outcomes

### v1

@@ D.3 Larger sample sizes
Here we present results from the same data generating process as the distributed lag simulation study in Section 5.1, however, we increase the sample size to $n=1500$. The results can be found across Figures D11 - D15. The results are largely similar to what was seen for smaller sample sizes, with a few expected, but notable, changes. Overall, the results point to the same finding that our proposed DLNM estimator generally outperforms estimators that do not perform clustering, and our approach without dimension reduction, while computationally slower, outperforms the one with dimension reduction as expected. One key difference in these results is that the clustering for the $\bm{\omega}_{kj}$ terms is significantly improved relative to the smaller sample size setting. This can be seen in Figure D12 where there is much clearer clustering structure, or in Figure D14 where there are fewer data sets for which the estimated distributed lag curves are placed in the wrong cluster. This leads to improved credible interval coverages, as seen in Figure D13. We now see that it is also preferable (in terms of MSE) to use our DLNM model with clustering across both exposures and outcomes for estimating $f_{kj}(\cdot)$, as seen in Figure D15. As a general rule, the $\bm{\omega}_{kj}$ parameters are only weakly identified, and are very difficult to estimate from small sample sizes, because of the high degree of correlation in the exposures across time. These issues improve as the sample size increases, which leads to the improved clustering performance, and the improved results when performing inference on the $\bm{\omega}_{kj}$ parameters.
@@
Figure D11: Mean squared error for estimating $\bm{\omega}_{kj}$ across all estimators when the sample size is $n=1500$. MSE is scaled so that the DLNM estimator has an MSE of 1 in all scenarios.
@@
Figure D12: Posterior probabilities of pairwise clustering for both sets of model parameters, averaged across simulated datasets when the sample size is $n=1500$.
@@
Figure D13: 90% credible interval coverage for estimating $\bm{\omega}_{kj}$ across all estimators when the sample size is $n=1500$.
@@
Figure D14: Estimates of $\bm{\omega}_{kj}$ for the DLNM estimator when the sample size is $n=1500$. Individual grey lines correspond to individual data sets, while the red line represents the truth, and the black line represents the mean across data sets.
@@
Figure D15: Mean squared error for estimating $f_{kj}(\cdot)$ across all estimators when the sample size is $n=1500$. MSE is scaled so that the DLNM estimator has an MSE of 1 in all scenarios.
Appendix E Additional NMMAPS Analysis Results
In Figure E16 we report mean outcome differences corresponding to an increase in ozone from 1SD below its mean to 1SD above at each lag $l$, holding exposures at all other lags constant. Results without clustering are reported on the left hand column; results with clustering are reported on the right. Overall, evidence of associations with ozone was low. Both approaches yielded constant null effects of ozone on cardiovascular deaths. Associations with respiratory deaths were strongest at lag 0 across both approaches; without clustering there was a slight increase in associations at lag 14, whereas with clustering lag effects decreased monotonically with lag. In both cases estimated associations were dwarfed by uncertainty. For other (non-accidental) deaths, the non-clustering approach found small protective effects of ozone at lag 0, although there was high uncertainty. The clustering approach reduced this coutnerintuitive result, however; associations were estimated to be constant and null across all lags.
@@ Appendix E Additional NMMAPS Analysis Results
Figure E16: Mean difference in outcomes increasing O3 from 1SD below the mean to 1SD above at each lag (holding all else constant). Three rows correspond to (log) cardiovascular, respiratory and other (non-accidental) deaths. Left column is with clustering; right column without clustering.
Estimated lag weight functions for PM10 and O3 are reported in Figures E17 E18, respectively. For identifiability, weights at lag 0 were constrained to be positive. (Note that the apparent ‘reflection’ in the top row of E17 is simply an artifact of the identifiability constraint, because lag-0 weights were estimated to be close to 0.) Overall results were similar to those above. For the non-clustered approach, the PM10-cardiovascular association exhibited U-shaped lag profiles, with strongest weights at days 7–9. The O3-cardiovascular weights were effectively constant across lags, and other lags were roughly linear. Clustering appeared to have the largest effect on the PM10-cardiovascular association, with lag profiles being a combination of U-shaped functions and linear functions. The O3-cardiovascular weights remained largely constant across lags. Other exposure-outcome pairs again had linear lag profiles, although weights at lag 13 were closer to null than in the non-clustering approach.
Figure E17: Lagged weight functions for PM10 analysis, constrained to have positive weight at lag 0. Three rows correspond to (log) cardiovascular, respiratory and other (non-accidental) deaths. Left column is without clustering; right column is with clustering.
Figure E18: Lagged weight functions for O3 analysis, constrained to have positive weight at lag 0. Three rows correspond to (log) cardiovascular, respiratory and other (non-accidental) deaths. Left column is with clustering; right column without clustering.

### v2

@@ G.3 Larger sample sizes
Here we present results from the same data generating process as the distributed lag simulation study in Section 5.1, however, we increase the sample size to $n=1500$. The results can be found across Figures G11 - G15. The results are largely similar to what was seen for smaller sample sizes, with a few expected, but notable, changes. Overall, the results point to the same finding that our proposed DLNM estimator (or DLNM with dimension reduction) generally outperforms estimators that do not perform clustering. One key difference in these results is that the clustering for the $\bm{\omega}_{kj}$ terms is improved relative to the smaller sample size setting. This can be seen in Figure G12 where there is clearer clustering structure, or in Figure G14 where there are fewer data sets for which the estimated distributed lag curves are placed in the wrong cluster. This leads to improved credible interval coverages, as seen in Figure G13. We also see that it is preferable (in terms of MSE) to use our DLNM model with clustering across both exposures and outcomes for estimating $f_{kj}(\cdot)$, as seen in Figure G15. As a general rule, the $\bm{\omega}_{kj}$ parameters are only weakly identified, and are very difficult to estimate from small sample sizes, because of the high degree of correlation in the exposures across time. These issues improve as the sample size increases, which leads to the improved clustering performance, and the improved results when performing inference on the $\bm{\omega}_{kj}$ parameters.
@@
Figure G11: Mean squared error for estimating $\bm{\omega}_{kj}$ across all estimators when the sample size is $n=1500$. MSE is scaled so that the DLNM estimator has an MSE of 1 in all scenarios.
@@
Figure G12: Posterior probabilities of pairwise clustering for both sets of model parameters, averaged across simulated datasets when the sample size is $n=1500$.
@@
Figure G13: 95% credible interval coverage for estimating $\bm{\omega}_{kj}$ across all estimators when the sample size is $n=1500$.
@@
Figure G14: Estimates of $\bm{\omega}_{kj}$ for the DLNM estimator when the sample size is $n=1500$. Individual grey lines correspond to individual data sets, while the red line represents the truth, and the black line represents the mean across data sets.
@@
Figure G15: Mean squared error for estimating $f_{kj}(\cdot)$ across all estimators when the sample size is $n=1500$. MSE is scaled so that the DLNM estimator has an MSE of 1 in all scenarios.
G.4 Sensitivity to choice of $C$
Here we assess the robustness of our approach to the choice of $C$, the number of mixtures within our finite mixture model. If one is considered about ensuring a large enough value of $C$ is chosen, then we recommend setting $C=K\times P$, which is the maximum number of distinct exposure-response functions that are needed, even when there is no clustering across exposures or outcomes. Ideally, if $C$ is chosen to be sufficiently large to ensure that enough distinct groups are available, then increasing $C$ should have little bearing on the final results as the model’s clustering properties should remove the effects of any unnecessary groups in the finite mixture model. To assess this empirically, we run the simulations from Section 5.1 in the manuscript, though we vary the choice of $C\in\{10,20\}$. In this setting, we have $K=4$ and $P=5$, so 20 clusters is the maximum possible number of distinct clusters. In the simulation study in the manuscript, we fixed $C=10$ throughout, and here we evaluate the sensitivity to this choice. The results can be found in Figures G16, G17, and G18. We see that the choice of $C$ does not make a large impact on the overall results. Both models have nearly identical 95% coverage rates across all exposures and outcomes examined, and the mean squared error for either estimating $\bm{\omega}_{kj}$ or $f_{kj}(\cdot)$ are similar between the two results. The MSE is slightly higher generally for $C=20$, though there are cases where it is higher when $C=10$ and the differences between the two are rather small. This is expected due to the clustering properties of the model that ensure unnecessary clusters do not have a large impact on the resulting estimates. More problematic would be to consider values of $C$ that are too small, which would force parameters to be clustered together that otherwise would not be, though this is easy to avoid by setting $C$ to be a relatively large value, where relative is based on the value of $K\times P$.
Figure G16: Mean squared error for estimating $\bm{\omega}_{kj}$ across all estimators using different values of $C$.
Figure G17: 95% credible interval coverage for estimating $\bm{\omega}_{kj}$ across all estimators using different values of $C$.
Figure G18: Mean squared error for estimating $f_{kj}(\cdot)$ across all estimators when using different values of $C$.
Appendix H Additional NMMAPS Analysis Results
Following chen2019distributed, we adjusted for: day of week (categorical), seasonality (cyclic basis splines for month, dim=4), year (thin plate splines, dim=6), and mean daily temperature and dew point temperature per day (thin plate splines, dim=6) and averaged over the prior three days (thin plate splines, dim=3), and we allowed different covariate associations for each outcome.
In Figure H19 we report results of a preliminary analysis using separate additive models based on single-day exposures. We found no statistically significant associations with same day exposure, but this ignores the fact that exposures may have delayed effects, hence motivating lagged exposure models like DLNMs.
Figure H19: Results of preliminary fits of separate models based on single-day exposures. Three rows correspond to (log) cardiovascular, respiratory and other (non-accidental) deaths. Left column corresponds to associations with same-day PM10; right is same-day O3.
In Figure H20 we report results of a preliminary analysis using separate lagged WQS models for each outcome (using quintiles, and $b=100$ bootstrap resamples). Results suggest weak negative associations with mixture, but uncertainty in time-varying exposure weights are ignored.
Figure H20: Results of preliminary fits of separate models via lagged WQS. Three rows correspond to (log) cardiovascular, respiratory and other (non-accidental) deaths.
In Figure H21 we report pairwise posterior clustering probabilities for the proposed MV-DLNM.
Figure H21: Pairwise posterior clustering probabilities for Chicago analysis. Left plot is for $\bm{\beta}$; right plot for $\bm{\theta}$
In Figures H22 and H23 we report mean outcome differences corresponding to an increase in each exposure from 1SD below their mean to 1SD above at each lag $l$, holding exposures at all other lags constant. Results for separate models are reported on the left hand column; MV-DLNM without clustering in the middle; results with clustering on the right.
@@ Appendix H Additional NMMAPS Analysis Results
Figure H22: Mean difference in outcomes increasing PM10 from 1SD below the mean to 1SD above at each lag (holding all else constant). Three rows correspond to (log) cardiovascular, respiratory and other (non-accidental) deaths. Left column is with separate model fits; middle is MV-DLNM without clustering; right is MV-DLNM with cluster-inducing priors.
Figure H23: Mean difference in outcomes increasing O3 from 1SD below the mean to 1SD above at each lag (holding all else constant). Three rows correspond to (log) cardiovascular, respiratory and other (non-accidental) deaths. Left column is with separate model fits; middle is MV-DLNM without clustering; right is MV-DLNM with cluster-inducing priors.
In Table H1 we quantify relative widths of credible intervals, comparing both the separate model fits and the MV-DLNM without clustering to the MV-DLNM with cluster-inducing priors.
Table H1: Relative interval width for lagged contrasts vs. proposed approach. “Sep.” represents separate model fits for each outcome; “MV-DLNM (no clustering)” represents the multivariate model but without cluster-inducing priors. Values greater than 1 indicate wider intervals than the proposed approach
Exposure Outcome Sep. MV-DLNM (no clustering)
PM10 Cardio. 3.60 3.27
Resp. 4.68 5.29
Other 2.13 2.02
O3 Cardio. 1.39 1.41
Resp. 1.42 1.54
Other 2.33 2.12
Average 2.59 2.61
Appendix I Alternative Model Specification: Non-Separable DLNM
We consider here the DLNM specification of gasparrini2017penalized:
$y_{ik}$ $=g_{k1}({\mathbf{x}}_{i1})+\cdots+g_{kP}({\mathbf{x}}_{iP})+\bm{z}_{i}^{T}\bm{\beta}_{Zk}+\xi\sigma_{k}u_{i}+\epsilon_{ik},~~k=1,\dots,K$
$\approx\sum_{l=1}^{L}h_{k1}({{x}}_{i1l},l)+\cdots+\sum_{l=1}^{L}h_{kP}({{x}}_{iPl},l)+\bm{z}_{i}^{T}\bm{\beta}_{Zk}+\xi\sigma_{k}u_{i}+\epsilon_{ik}$
where $h_{kj}(x,l)$ is a bi-dimensional dose-lag-response function (gasparrini2017penalized).
We then adopt a tensor product basis representation, which requires specifying marginal bases for the $x$- and time-dimensions. We use the same bases as described in the main text. Applying these (known) basis functions to $\bm{x}_{ij}$ and $1,\dots,L]^{T}$ yields corresponding basis matrices $\mathbf{R}_{ij}$ (of dimension $L\times d$) and $\bm{\Psi}$ (of dimension $L\times m$). This yields the representation:
$y_{ik}$ $\approx\beta_{k0}+\mathbf{b}_{i1}^{T}\bm{\beta}_{k1}+\cdots+\mathbf{b}_{iP}^{T}\bm{\beta}_{kP}+\bm{z}_{i}^{T}\bm{\beta}_{Zk}+\xi\sigma_{k}u_{i}+\epsilon_{ik}$
where
$\mathbf{b}_{ij}^{T}$ $=\mathbf{1}^{T}_{L}\left[\left(\mathbf{R}_{ij}\otimes\mathbf{1}_{m}^{\top}\right)\odot\left(\mathbf{1}_{d}^{\top}\otimes\bm{\Psi}\right)\right],$
is a known vector, and $\bm{\beta}_{kj}$ is an unknown coefficient vector of length $d\cdot m$. Note as in the main text we have included an intercept and imposed centering constraints to the basis representations (wood2017generalized).
To encourage smoothness in the bidimensional function we can adopt Gaussian priors for $\bm{\beta}_{kj}$ with mean $\mathbf{0}$ and precision matrix $\lambda^{\beta}\widetilde{\mathbf{P}}$ where $\widetilde{\mathbf{P}}=(\bm{\Sigma}_{0}^{-1}\otimes\mathbf{I}_{m})+(\mathbf{I}_{d}\otimes\bm{\Sigma}_{\theta}^{-1})$, where $\bm{\Sigma}_{0}^{-1}$ and $\bm{\Sigma}_{\theta}^{-1}$ are penalty matrices as in the main text, and $\lambda^{\beta}$ and is a smooth penalty parameters defined in the main text. (Here we use a single penalty parameter for simplicity, but this can be extended to more complex penalties for different dimensions; sampling follows as in bach2025anisotropic.)
I.1 Proposed Cluster-Inducing Priors
As in the main model specification, we aim to borrow strength across exposures and outcomes. A key departure for this model specification, however, is that there is no $\bm{\theta}_{kj}$; all parameters related to the dose-lag-response function for the $k^{th}$ outcome and the $j^{th}$ exposure is captured by the single (albeit much larger) vector $\bm{\beta}_{kj}$. An advantage of this non-separable specification is that we don’t need to propose a dependent co-clustering strategy; we can simply place cluster-inducing priors on the larger coefficient vectors $\bm{\beta}_{kj}$. A disadvantage, however, is this doesn’t allow separate levels of clustering lag profiles and exposure response relationships. Specifically we propose DP mixture priors $\bm{\beta}_{kj}\sim G_{\beta}$ where $G_{\beta}\sim DP(\alpha_{\beta},G_{0,\beta})$, and we use the same parameterization as in the main manuscript. As centering distribution $G_{0,\beta}$, we adopt the smoothness-inducing Gaussian described above. We use the same default priors on all other parameters.
I.2 Modified MCMC Sampler
  1. 1.
    For $k=1,\dots,K$, $j=1,\dots,P$, sample cluster IDs:
[…]
  2. 2.
    $\bm{\beta}^{*}_{c}|$... $\sim N(\bm{\mu}^{*},\bm{\Sigma}^{*})$
    $\bm{\mu}^{*}$ $=\bm{\Sigma}^{*}\times\left(\lambda^{\beta}\bm{\mu}_{0}^{T}\widetilde{\mathbf{P}}+\sum_{k}^{K}\sigma^{-2}_{k}\left[\sum_{j:Z^{\beta}_{kj}=c}\mathbf{B}_{j}^{T}\right]\left[\mathbf{y}_{k}-\beta_{0k}\mathbf{1}-\xi\sigma_{k}\mathbf{u}-\sum_{j:Z^{\beta}_{kj}\neq c}\mathbf{B}_{j}\bm{\beta}_{kj}\right]\right)$
    $\bm{\Sigma}^{*}$ $=\left(\lambda^{\beta}\widetilde{\mathbf{P}}+\sum_{k}^{K}\sigma^{-2}_{k}\left[\sum_{j:Z^{\beta}_{kj}=c}\mathbf{B}_{j}^{T}\right]\left[\sum_{j:Z^{\beta}_{kj}=c}\mathbf{B}_{j}\right]\right)^{-1}$
  3. 3.
    Sample penalty term $\lambda^{\beta}$ from full conditional:
    $((\lambda^{\beta})|\dots)$ $\propto Gamma(a_{\lambda}^{\beta}+\frac{C}{2}\text{rank}(\widetilde{\mathbf{P}}),~~b_{\lambda}^{\beta}+\frac{1}{2}\sum_{c=1}^{C}{\bm{\beta}_{c}^{*}}^{T}\widetilde{\mathbf{P}}\bm{\beta}_{c}^{*})$
All other updates—for $V^{\beta}_{c}$, $\alpha_{\beta}$, $u_{i}$, $\xi$, $\sigma^{2}_{k}$, $\beta_{0k}$ and $\bm{\beta}_{Zk}$—are the same as in Appendix D, replacing $\sum_{j=1}^{p}\mathbf{b}_{i,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}$ with $\sum_{j=1}^{p}\mathbf{b}_{ij}^{T}\bm{\beta}_{kj}$.
I.3 Non-Separable DLNM: Simulations
In this section we evaluate the proposed approach within the nonseparable distributed lag framework described above. We explore three distinct approaches that vary in their degree of clustering:
  1. 1.
    The proposed nonseparable formulation with clustering across exposures and outcomes (DLNM)
  2. 2.
    The proposed nonseparable formulation without any clustering (DLNM no clustering)
  3. 3.
    The proposed nonseparable formulation applied separately to each outcome, which allows for clustering across exposures within an outcome, but not across outcomes (DLNM separate)
As in the manuscript simulations, we set the sample size to be $n=500$, the number of time points to be $L=52$, the number of exposures to be $P=5$, and the number of outcomes to be $K=4$. The exposures have an autoregressive correlation structure that ensures exposures at successive time points have correlation 0.85, while the exposures are uncorrelated with each other at each point in time. Responses were generated from a multivariate normal distribution with mean:
$E[\bm{Y}\mid\bm{X}]$ $=\left[\begin{array}[]{l}g_{1}(\mathbf{x}_{i1})+g_{2}(\mathbf{x}_{i2})+g_{3}(\mathbf{x}_{i3})+g_{1}(\mathbf{x}_{i4})+g_{1}(\mathbf{x}_{i5})\\ g_{1}(\mathbf{x}_{i1})+g_{2}(\mathbf{x}_{i2})+g_{3}(\mathbf{x}_{i3})+g_{1}(\mathbf{x}_{i4})+g_{1}(\mathbf{x}_{i5})\\ g_{3}(\mathbf{x}_{i1})+g_{3}(\mathbf{x}_{i2})+g_{3}(\mathbf{x}_{i3})+g_{3}(\mathbf{x}_{i4})+g_{3}(\mathbf{x}_{i5})\\ g_{3}(\mathbf{x}_{i1})+g_{3}(\mathbf{x}_{i2})+g_{3}(\mathbf{x}_{i3})+g_{3}(\mathbf{x}_{i4})+g_{3}(\mathbf{x}_{i5})\\ \end{array}\right]$
where
$g_{m}(\mathbf{x}_{ij})=\sum_{\ell=1}^{L}f_{m}\cdot w_{m}\left(x_{ij\ell},\ell\right)$
, and
[…]
$f_{2}\cdot w_{2}\left(x_{ij\ell},\ell\right)={l}f_{2}\left(x_{ij\ell}\right)\cdot w_{2}(\ell)$
$f_{3}\cdot w_{3}\left(x_{ij\ell},\ell\right)=\left\{\begin{array}[]{l}f_{3}\left(x_{ij\ell}\right)\cdot w_{1}(\ell)\quad\text{ if }x\geq 0\\ f_{3}\left(x_{ij\ell}\right)\cdot w_{3}(\ell)\quad\text{ if }x<0\end{array},\right$.
where $w_{1}(\ell)$ is flat, $w_{2}(\ell)$ is decreasing, $w_{3}(\ell)$ is increasing, and:
$f_{1}(x)$ $=0.04\cdot(x-2)$
$f_{2}(x)$ $=\sum_{p=0}^{4}0.3\delta_{p}(x^{p}-5^{p})$
$f_{3}(x)$ $=0.4\left(\phi(x;\mu=1.5,\sigma=2)+\phi(x;\mu=7.5,\sigma=1)\right)$
$~~-0.4\left(\phi(5;\mu=1.5,\sigma=2)+\phi(5;\mu=7.5,\sigma=1)\right)$
where $\bm{\delta}=[0.2118881,0.1406585,-0.0982663,0.0153671,-0.0006265]^{T}$. Note that these functional forms here were motivated by those of gasparrini2017penalized. The results can be found in Figures I24, I25, and I26. Overall, we see that clustering dramatically improves performance within the non-separable model formulation as well. The model with clustering significantly outperforms the other two models in terms of mean squared error. We see that as in the manuscript, this can at times lower the 95% interval coverage rate due to bias incurred from clustering, but this comes with large reductions in MSE. Lastly, we can see from the posterior probabilities of clustering that exposure-outcome pairs with similar effects on the outcome are indeed clustered together with high probability, allowing for borrowing of information across exposures and outcomes.
Figure I24: Posterior probabilities of clustering for model parameters, averaged across simulated datasets in the non-separable setting.
Figure I25: 95% credible interval coverage for estimating $f_{kj}(\cdot)$ across all estimators in the non-separable setting.
Figure I26: Mean squared error for estimating $f_{kj}(\cdot)$ across all estimators in the non-separable setting.
Appendix J Alternative Model Specification: Non-Additivity
Here we discuss extensions of our approach to allow for non-additivity, with a particular focus on the distributed lag application of our general modeling approach. We do not discuss non-additivity in the adaptive multiple index model setting described in Section 4, because the effects of the exposures in that setting are already allowed to be non-additive. We focus on the distributed lag nonlinear model described in Section 3, as our current approach does not allow for the effects of one exposure to depend on the values of the others. In this setting, interactions between exposures can occur in complicated ways as exposure $j$ at time point $t$ can interact with exposure $j^{\prime}$ at time point $t^{\prime}$. To allow for arbitrary interactions such as this, we can fit the proposed extension of our model given by:
$y_{ik}$ $=\sum_{j=1}^{P}f_{kj}({\mathbf{x}}_{ij}^{T}{\bm{\theta}}_{kj})+\sum_{j=1}^{P}\sum_{j^{\prime}=j+1}^{P}\tilde{f}_{kjj^{\prime}}({\tilde{\mathbf{x}}}_{ijj^{\prime}}^{T}{\tilde{\bm{\theta}}}_{kjj^{\prime}})+\bm{z}_{i}^{T}\bm{\beta}_{Zk}+\xi\sigma_{k}u_{i}+\epsilon_{ik},$
where we now have that ${\tilde{\mathbf{x}}}_{ijj^{\prime}}$ contains all of the $L^{2}$ values for exposures $j$ and $j^{\prime}$ at all possible pairs of time points, and ${\tilde{\bm{\theta}}}_{kjj^{\prime}}$ is the corresponding parameter vector. The parameters ${\tilde{\bm{\theta}}}_{kjj^{\prime}}$ can be viewed as a (vectorized) two-dimensional matrix of dimension $L\times L$ that corresponds to all possible pairs of interactions between exposures $j$ and $j^{\prime}$. The same co-clustering ideas developed in the manuscript for $f_{kj}(\cdot)$ and $\bm{\theta}_{kj}$ could be applied in a similar manner for $\tilde{f}_{kjj^{\prime}}(\cdot)$ and ${\tilde{\bm{\theta}}}_{kjj^{\prime}}$. The main difficulty for implementation of this approach is the high-dimensional nature of the parameter vector as we would have $L^{2}$ parameters for each pair of exposures, and there are $\binom{P}{2}$ such pairs. One option is to do as we did for the $\bm{\theta}_{kj}$ parameters and assume smoothness over time, which can be incorporated through appropriate prior distribution variances for these parameters. A similar approach could be taken for ${\tilde{\bm{\theta}}}_{kjj^{\prime}}$ although it is not trivial to develop a penalty matrix that incorporates smoothness in two dimensions, though this could be a topic for future research. Alternative approaches would be to impose smoothness by approximating ${\tilde{\bm{\theta}}}_{kjj^{\prime}}$ as a linear combination of two-dimensional basis functions as done in antonelli2024multiple, though this requires specifying basis functions a priori and may not work well if the true coefficient matrix is not well captured by the chosen basis functions. This can be viewed as a special case of the dimension reduction that we applied in the manuscript, and the same co-clustering ideas could apply immediately. In practice, however, one should be careful implementing such a model because interaction effects in distributed lag settings for environmental mixtures are likely very small and difficult to detect with even moderate sample sizes. Additionally, increasing the complexity of the model unnecessarily may lead to less stable and highly variable results unless appropriate regularization or variable selection is implemented. However, such regularization or variable selection will likely remove any interaction effects, even if they exist, due to their small signal.
Appendix K Identifiability & Orthogonalization
As in a typical mixed effects model, the fixed effects are of primary interest, and the random effects serve to model residual correlation that remains after adjusting for exposure/covariate fixed effects. The clustering strategy aims to borrow strength in those exposure effects across different outcomes, whereas inclusion of a random effect $u_{i}$ serves to absorb residual correlation not captured by between-subject variation in exposure levels. In our case, however, the fact that the exposure effects are encouraged by the cluster-inducing priors to be similar means there could be an identifiability problem. This is analogous to a phenomenon known as spatial confounding (not to be conflated with the causal notion of confounding), wherein near non-identifiability can arise when spatially varying covariates are included in a model in addition to spatially varying random effects. In the spatial statistics literature, several methods have been proposed to solve this issue, including restricted spatial regression (RSR; reich2006effects). The RSR approach reparameterizes the model such that the random effects are orthogonal to the column space of the fixed effects. For example replacing the vector of random effects $\bm{u}$ in the model
$\mathbf{y}_{\cdot k}$ $\approx\beta_{k0}\mathbf{1}+\mathbf{B}_{\cdot k}\bm{\beta}_{k\cdot}+\bm{Z}\bm{\beta}_{Zk}+\xi\sigma_{k}\bm{u}+\bm{\epsilon}_{\cdot k}$
$\bm{u}$ ${\sim}N(\mathbf{0},\mathbf{I})$
with $\bm{u}^{*}=[\mathbf{I}-\mathbf{P}]\bm{u}$, where $\mathbf{P}$ is the usual projection matrix onto the column space of $\widetilde{\mathbf{B}}:=[\mathbf{1}|\mathbf{B}|\mathbf{Z}]$. A computationally efficient approach for implementing the RSR approach is conditional kriging (hanks2015restricted; rue2005gaussian), wherein the MCMC sampler draws random effects from the usual un-constrained posterior $\bm{u}\sim N(\bm{\mu}_{u},\bm{\Sigma}_{u})$ and then orthogonalizes them: $\bm{u}^{*}=\bm{u}-\bm{\Sigma}_{u}\widetilde{\mathbf{B}}[\widetilde{\mathbf{B}}^{T}\bm{\Sigma}_{u}\widetilde{\mathbf{B}}]^{-1}\widetilde{\mathbf{B}}\bm{b}$.
A challenge in our setting is that the equivalent column space here depends on unknown parameters that are being sampled, $\bm{\theta}_{kj}$, so MCMC becomes very challenging. As an approximation we simply change the $u$ update to be orthogonal to $\mathbf{B}_{\cdot,\bm{\theta}_{1,\theta}}$ (but under the non-separable parameterization of I the exact update is straightforward).
K.1 Identifiability Simulations
Here we assess the potential impact of the identifiability issue described above, and the extent to which the conditioning by kriging approximation is able to address this issue. Throughout these simulations we focus on the version of our approach that implements clustering across both exposures and outcomes, and does incorporate dimension reduction when estimating $\bm{\theta}_{kj}$. We explore our approach both with and without conditioning by kriging, and we evaluate its performance in two simulation scenarios. The first simulation scenario is identical to the distributed lag scenario explored in the manuscript, where there should be no identifiability issues due to the fact that the exposure effects vary across the outcomes. The second simulation is identical to the first simulation with respect to the distribution of the exposures, though it varies with respect to the relationship between the exposures and the outcome. It represents the most extreme scenario where the effect of each exposure on each of the outcomes is identical, both with respect to the distributed lag parameters $\bm{\theta}_{kj}$ and the exposure response function that is captured by the $\bm{\beta}_{kj}$ parameters. Specifically, we generate the outcome from a multivariate normal distribution with mean:
$E[\bm{Y}\mid\bm{X}]$ $=\left[\begin{array}[]{l}f_{2}\left(\bm{x}_{1}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{3}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{4}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{5}^{T}\bm{\omega}_{2}\right)\\ f_{2}\left(\bm{x}_{1}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{3}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{4}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{5}^{T}\bm{\omega}_{2}\right)\\ f_{2}\left(\bm{x}_{1}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{3}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{4}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{5}^{T}\bm{\omega}_{2}\right)\\ f_{2}\left(\bm{x}_{1}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{3}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{4}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{5}^{T}\bm{\omega}_{2}\right)\end{array}\right],$
where, as in the manuscript, we have that $f_{2}(a)=0.24(0.3a)^{2}$ and $\bm{\omega}_{2}$ is the same increasing function used in the manuscript.
The results can be seen in Figure K27, where we see that the decision about whether to orthogonalize the random effects does not substantially change the results. In both scenarios, the mean squared error for estimating the weight parameters or the entire mean function are largely similar, though slightly worse when implementing conditioning by kriging, and the coverage rates are quite similar between the two approaches in both scenarios. This is expected in scenario 1 where we should not have any identifiability issues, but is somewhat surprising for scenario 2 where it is expected that we would have an identifiability issue between the magnitude of the random effect variance and the magnitude of the exposure effects. While this identifiability issue could manifest in principle, it is not one that we encounter even in this extreme scenario. The reason for this is likely that it takes a reasonable number of MCMC samples for the clustering indicators to center on their correct values, which in scenario 2 would have all exposure and exposure-response parameters to be clustered together. Identifiability only becomes an issue when the exposure effects are identical across all outcomes, and it appears that the estimate of the random effect variance centers around the correct value in the early MCMC samples while the exposure effects are still not identical, which leads to the performance seen here. Overall this shows that identifiability is likely not a huge concern for our approach, particularly because it only is an issue in extreme scenarios.
Figure K27: Comparison of results with and without orthogonalizing the random effects.